# Quantum dissipative dynamics of driven Duffing oscillator near attractors

### v1

@@ III.1.3 Effective temperature
Next, we calculate the stationary distribution over the levels of the HAS and the effective temperature near the bottom. It is important to note that the annihilation operator $a$ is for the Fock state $|N\rangle$, which decreases the Fock state from a higher level to the next lower level $a|N\rangle=\sqrt{N}|N-1\rangle$. However, in our case, the eigenstate of quasienergy $|n\rangle$ is the superposition of the Fock state $|N\rangle$. As a result, the annihilation operator $a$ can either decrease or increase the state $|n\rangle$ even at zero temperature.
Under the assumption of weak damping ($\kappa\ll E_{n}-E_{n+1}$), the off-diagonal matrix elements on the state $|n\rangle$ are very small. Thus, we can only keep the diagonal elements. Here, we assume that the stationary density matrix is diagonal and denote the diagonal terms $p_{n^{\prime}}=\langle n^{\prime}|\rho|n^{\prime}\rangle$. The master equation (II.2) can be simplified into a balance equation Dyk06
@@
for level $|n\rangle$. Fig. 6 illustrates how $N_{eff}$ varies from the lowest to higher levels. The zero-order term yields a constant effective temperature. When we include the correction to the order of $\lambda$, the correction leads to changes in $N_{eff}$ showing a good agreement with the numerical results for levels near the bottom.
@@
Figure 8: Renormalized effective Bose distribution $\tilde{N}$ due to dephasing effects for (a) the LAS and (b) the HAS. We compare the theoretical predictions from Eq. (36) with the exact numerical results and extract the linear relationships.
@@ III.2 Strong Damping and Dephasing
In this section, we explore the dynamics and stationary state of the system under conditions of strong damping and dephasing, which can significantly alter the behavior predicted by the weak damping approximation.
@@ III.2.1 Strong damping
In the regime of strong damping, the harmonic approximation, which leads to the level spacing $\Delta E_{n}$ that is independent of damping, is no longer valid. Instead, the level spacing undergoes slight renormalization for strong damping. Such effect can be observed through the emission spectrum $S(\omega)$, representing the spectral density of photons emitted by the driven resonator, is given by
@@
Our method inherently incorporates damping effects into the calculation of $\Delta E_{n}$. For finite damping, the squeezing parameters $u$ and $v$ are complex numbers, determined by the steady-state conditions given in Eq. (29). Substituting them into Eq. (26), we obtain the effective frequency $\omega_{eff}=\Delta E_{1}$. In Fig. 7, we compare the results obtained through the emission spectrum Eq. (34) with those from our perturbation theory, demonstrating satisfactory consistency. The minor discrepancy arises primarily from higher-level spacings $\Delta E_{n}(n>1)$, which are typically smaller than the first-level spacing $\Delta E_{1}$. For more accurate results, the average of all level spacings should be considered in the calculations.
@@ III.2.2 Dephasing
To incorporate dephasing, we introduce the dephasing term $\eta^{ph}{\cal D}[a^{\dagger}a]{\rho}$ into the master equation (7). For convenience, we define a generalized Lindblad operator ${\cal L}[A;B]\rho\equiv 2A\rho B-BA\rho-\rho AB$. In the spirit of the rotating wave approximation, we obtain the renormalized master equation for the displaced and squeezed density ioperator $\bar{\rho}=S^{\dagger}D^{\dagger}{\rho}DS$ (see detailed derivation in Appendix B)
@@
We verify our predictions for renormalized Bose distribution by comparing them with exact numerical simulations according to the probabilities over the two lowest levels, specifically $\tilde{N}=p_{2}/(p_{1}-p_{2})$. In Fig. 8, we plot and extract the renormalized Bose distribution as a function of the dephasing, which demonstrates an excellent agreement between the numerical results and the analytical predictions given by Eq. (36) for both LAS and HAS.
@@ IV Conclusions
In this work, we have investigated the quantum dissipative dynamics of a driven Duffing oscillator near the bottoms of its stable states. We elucidated the intricate interplay between the nonlinearity, quantum fluctuations, and the influence of an external driving field. We formulated an effective quantum master equation that encompasses quantum and thermal fluctuations, strong damping, and dephasing within a unified framework. We have developed a refined perturbation theory to analyze the quantum dynamics near both the LAS and the HAS of the Duffing oscillator. While the LAS behavior can be approximated using a harmonic oscillator model, the HAS exhibits more complex behavior due to the significant nonlinear terms. Due to quantum fluctuations, even at zero temperature, higher energy levels near the bottom of the potential well are excited. We calculated the level spacing and effective temperature near the bottom of the HAS and compare them with numerical simulations, demonstrating the accuracy and utility of our proposed approach.
@@ ACKNOWLEDGMENTS
This work was supported by Natural Science Foundation of China (Grant No. 12475025).
@@ Appendix A Double Perturbation Theory
which gives us the second order perturbative result:

### v2

@@ III.1.3 Effective temperature
Next, we calculate the stationary distribution over the levels of the HAS and the effective temperature near the bottom. It is important to note that the annihilation operator $a$ is for the Fock state $|N\rangle$, which decreases the Fock state from a higher level to the next lower level $a|N\rangle=\sqrt{N}|N-1\rangle$. However, in our case, the eigenstate of quasienergy $|n\rangle$ is the superposition of Fock states $|N\rangle$. As a result, the annihilation operator $a$ can either decrease or increase the state $|n\rangle$ even at zero temperature.
Under the assumption of weak damping ($\kappa\ll E_{n}-E_{n+1}$), the off-diagonal matrix elements on the state $|n\rangle$ are very small. Thus, we can only keep the diagonal elements. Here, we assume that the stationary density matrix is diagonal and denote the diagonal terms as $p_{n^{\prime}}=\langle n^{\prime}|\rho|n^{\prime}\rangle$. The master equation (II.2) can be simplified into a balance equation Dyk06
@@
for level $|n\rangle$. Fig. 6 illustrates how $N_{eff}$ varies from the lowest to higher levels. The zero-order term yields a constant effective temperature. When we include the correction up to the order of $\lambda$, the correction leads to changes in $N_{eff}$ showing a good agreement with the numerical results for levels near the bottom.
@@
Figure 8: Renormalized effective Bose distribution $\tilde{N}$ due to dephasing effects for (a) the LAS and (b) the HAS. We compare theoretical predictions from Eq. (36) with the exact numerical results and extract the linear relationships.
@@ III.2 Strong Damping and Dephasing
In this section, we explore the dynamics and the stationary state of the system under conditions of strong damping and dephasing, which can significantly alter the behavior predicted by the weak damping approximation.
@@ III.2.1 Strong damping
In the regime of strong damping, the harmonic approximation, which leads to the level spacing $\Delta E_{n}$ that is independent of damping, is no longer valid. Instead, the level spacing undergoes slight renormalization for strong damping. Such an effect can be observed through the emission spectrum $S(\omega)$, which represents the spectral density of photons emitted by the driven resonator and is given by
@@
Our method inherently incorporates damping effects into the calculation of $\Delta E_{n}$. For finite damping, the squeezing parameters $u$ and $v$ are complex numbers, determined by the steady-state conditions given in Eq. (29). Substituting these parameters into Eq. (26), we obtain the effective frequency $\omega_{eff}=\Delta E_{1}$. In Fig. 7, we compare the results obtained through the emission spectrum Eq. (34) with those from our perturbation theory, demonstrating satisfactory consistency. The minor discrepancy arises primarily from higher-level spacings $\Delta E_{n}(n>1)$, which are generally smaller than the first-level spacing $\Delta E_{1}$. For more accurate results, the average of all level spacings should be considered in the calculations.
@@ III.2.2 Dephasing
To incorporate dephasing, we introduce the dephasing term $\eta^{ph}{\cal D}[a^{\dagger}a]{\rho}$ into the master equation (7). For convenience, we define a generalized Lindblad operator ${\cal L}[A;B]\rho\equiv 2A\rho B-BA\rho-\rho AB$. In the spirit of the rotating wave approximation, we obtain the renormalized master equation for the displaced and squeezed density operator $\bar{\rho}=S^{\dagger}D^{\dagger}{\rho}DS$ (see the detailed derivation in Appendix B)
@@
We verify our predictions for the renormalized Bose distribution by comparing them with exact numerical simulations according to the probabilities over the two lowest levels, specifically, $\tilde{N}=p_{2}/(p_{1}-p_{2})$. In Fig. 8, we plot and extract the renormalized Bose distribution as a function of dephasing, which demonstrates an excellent agreement between the numerical results and the analytical predictions given by Eq. (36) for both LAS and HAS.
@@ IV Conclusions
In this work, we have investigated the quantum dissipative dynamics of a driven Duffing oscillator near the bottoms of its stable states. We elucidated the intricate interplay among the nonlinearity, quantum fluctuations, and the influence of an external driving field. We formulated an effective quantum master equation that encompasses quantum and thermal fluctuations, strong damping, and dephasing within a unified framework. We have developed a refined perturbation approach to analyze the quantum dynamics near both the LAS and the HAS of the Duffing oscillator. While the LAS behavior can be approximated using a harmonic oscillator model, the HAS exhibits more complex behavior owing to significant nonlinear terms. Because of quantum fluctuations, even at zero temperature, higher energy levels near the bottom of the potential well are excited. We calculated the level spacing and effective temperature near the bottom of the HAS and compared them with numerical simulations, demonstrating the accuracy and utility of our proposed approach.
@@ ACKNOWLEDGMENTS
This work was supported by the Natural Science Foundation of China (Grant No. 12475025).
@@ Appendix A Double Perturbation Theory
which gives us the second-order perturbative result: